[clas_members] RC daily summary for Saturday February 25

silvia at jlab.org silvia at jlab.org
Sat Feb 25 11:40:11 EST 2023

Day shift (Friday):
Beam went off before 9AM, work in the accelerator and in Hall-D to fix
their dipole magnet.
Target group removed the CH2 and replaced with ND3. Repolarized, faster
and reaching higher values than expected.
Work done on the CND (removed CAEN CFD, fixed one loose splitter cable).
Dan Carman measured helium poisoning on CTOF. Work done by Sergei.

Swing shift (Friday):
No beam for the whole shift.

Owl Shift:
Beam came back at 6AM, the plan is to take a Moeller run. Beam tuning,
harp scans look good but then MCC has troubles providing good positions
for Moeller run.

Day shift: beam tuning completed at 8:30 or so, but tagger magnet trips
and needs to be turned back on. Moeller run started at 9:25, ended at
9:50, beam polarization: 77.972 +/- 1.464. Tagger magnet is degaussed,
sent beam to FC, checked positions, turned on the detectors, RICH
recovery, recovered DAQ, and started taking production data on ND3 at
The target polarization reached a never-seen-before -51% right before
starting to get beam. At -49% now after 1 hour of beam.

The plan for today is to continue data taking on ND3, HWP IN. We will
change HWP sometime tomorrow, depending on how much beam we get.

No RC meeting today, enjoy your WE.

More information about the clas_members mailing list